Subject: APO 3.5-m Users Committee minutes 2/14/05
From: Bruce Gillespie
Submitted: Fri, 18 Feb 2005 10:14:19 -0700
Message number: 892
(previous: 891,
next: 893
up: Index)
APO 3.5-m Users Committee Phonecon, 2/14/05
Attending: Suzanne Hawley, Michael Strauss, Karl Glazebrook, Bruce
Balick, Bruce Gillespie, Rene Walterbos, Al Harper, Jon Holtzman
Absent: Don York, Russet McMillan, Fred Hearty, John Bally
Minutes taken by Bruce Gillespie
**********************************
NIC-FPS status:
Since Fred Hearty was in transit to APO and unavailable for this
teleconference, the NIC-FPS status and plans report was mainly given by
Suzanne and Bruce G. Bruce Balick asked about the availability of
NIC-FPS user documentation. Bruce G. said that there was a message
posted to apo35general from John Barentine that points to NIC-FPS
documentation (see message 851 in the apo35-general archive). It was
also pointed out that until all of the user documentation is in place,
Fred Hearty has offered to help NIC-FPS users personally, both before
and during their observing runs. Suzanne added that some of the ARC
institutions now have one (or more) "seasoned" NIC-FPS users, and that
new users could try consulting with them. She also pointed out that we
have decided to allow NIC-FPS be used as a primary or backup instrument
outside of the scheduled "campaign" schedule blocks, and that it will
be treated as a normally scheduled instrument in Q2 (sans etalon).
Also, Q2 will be the last quarter that GRIM2 is scheduled, so that
users who need to transition their science programs from GRIM2 to
NIC-FPS will need to plan according. There was some discussion of what
to do with GRIM2 once it's retired. Since it is ARC property, one
thought was to surplus the instrument to an ARC institution if anybody
wanted it. Al Harper expressed possible interest for UC.
The Fabry-Perot side of NIC-FPS is presently not working well enough
for science, and will be worked on during the next several weeks. If
this is unsuccessful, the F-P will be removed from the instrument and
is not likely to be reinstalled until after the summer shutdowns. Also,
the vacuum-hold problem we are presently experiencing with the
instrument is going to be worked on during the summer, most probably in
Boulder.
**********************************
Gamma-Ray Burst programs:
The GRB science document, being assembled by Don Lamb et al., will soon
be available. Suzanne will pass it on to the users committee members
and asked them to study it, share it with interested ARC astronomers,
and give her feedback. It was mentioned that the forthcoming document
should be treated as somewhat proprietary, i.e., don't post it in a
public place.
**********************************
Next Futures Committee meeting:
The next Futures Committee meeting will be held in Boulder on 3-4
March. At the last meeting in Chicago, discussions included the
UVirginia Triplespec instrument proposal. At the upcoming meeting,
there will be a presentation by Jian Ge on the Florida extra-solar
planet search proposal, and the committee will discuss the impacts and
potential interactions between the 2.5-m and 3.5-m telescopes at APO.
The Futures Committee will need to hear feedback from the community
regarding this proposal and its potential impacts. Suzanne asked that
the Users Committee members should talk to the Futures Committee
members at their respective institutions, in cases where it's not the
same person. She also mentioned that she had recently talked to George
Jacoby about possible future collaboration with WIYN. Suzanne will
discuss this with the Futures Committee next month. Michael added that
the SDSS project will be at the NSF review of the extension proposal
later this week. We are hoping to hear a decision from the NSF before
1 April. Rene said it would be useful to know the decision by the time
of the Futures Committee meeting next month, but that it would be
unlikely that we would hear anything by then. On another
futures-related topic, Suzanne said that she would talk to Karl at the
meeting about the JHU proposal for potential upgrades to DIS (see
below).
**********************************
Telescope & Instruments report:
Bruce G. reported that the telescope and instruments have been working
fairly well recently, although the weather has been unusually wet and
cloudy. Suzanne mentioned that the new top-end CDR is tentatively
scheduled for April, and that the new spare drive boxes are being
constructed in Seattle; both are planned for installation this year.
Bruce G. pointed out that with the uncertainty of SDSS-II, there has
been (and possibly will be more) staff attrition at the site that
affects both the 3.5-m and SDSS. Various contingency plans have, or
will be, put in place to minimize the consequences of reduced staffing
levels over the coming months, pending the outcome of the SDSS-II
proposal.
**********************************
DIS upgrade proposal:
Since DIS is our most-used instrument, some CIF funds have been
budgeted for enhancing its performance. Jon Holtzman is looking into
this, and has identified three relatively minor upgrades to improve
throughput:
o new gratings (to improve throughput, and chip and wavelength
coverage)
o UV throughput (fixing the optics to allow throughput below 3900
Angstroms)
o Dichroic split (to fix a dip in the throughput at the split between
red and blue)
Jon has asked the Instrument Development Group at JHU to make
recommendations on how to do these improvements, and it would be good
if Jon heard comments from DIS users about their grating parameter and
wavelength preferences. Also, Suzanne mentioned that the red-fringing
problem is being investigated, and wondered if anybody had solved it in
reductions. Michael said that you need to take flats for every science
exposure, but others doubted if this is enough. People with red data
who either have problems with fringing and/or have solved the fringing
problem are encouraged to contact their users committee representative
and Jon Holtzman. Rene mentioned that doing velocity work in the blue
is hard, because you need lots of light to get calibration lines, which
shift as the instrument rotates. Also, scattered light in the blue
side has been a problem for some. Jon said that one idea would be to
have all three gratings mounted at the same time, with the low-res
gratings mounted at a fixed angle in place of the imaging mirrors. He
wondered if this would be a problem, and most agreed that it wouldn't
because not many people ever use the imaging mirrors, and they could be
installed in advance if someone did need to use them. Also, if all
three turrets had gratings, doing slit-less spectroscopy would have the
same caveats.
**********************************
Other topics:
Rene asked for a summary of the summer shutdown plans. Suzanne said
that the plans are not exact yet, but we expect to have a one-week
shutdown in June to replace the telescope drives, and two-week
shutdowns in both July and August to do annual maintenance and install
the new top-end. All shutdowns are being targeted for bright time.
**********************************
Last month's minutes were approved without comment. Next phonecon will
be on Monday, March 28, at 8:30 AM Pacific Standard Time.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 892 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO